# algorithmic

### v1

@@ 3.2 Noise Level Correction
For $\text{dist}_{\mathcal{K}}(x_{t})$, calculating the ground-truth distance to the manifold is generally not feasible. Instead, we approximate it using the distance between the noisy sample and its clean counterpart in the forward diffusion process. Specifically, given the noisy samples $x_{t}$ generated by the diffusion process from $x_{0}\in\mathcal{K}$ in eq. 2, we estimate the distance as $\text{dist}_{\mathcal{K}}(x_{t})\approx|x_{t}-x_{0}|$. This approximation is reasonable because, according to the manifold hypothesis, the random noise $\epsilon$ is orthogonal to the manifold $\mathcal{K}$. Therefore, the projection of $x_{t}$ onto $\mathcal{K}$ satisfies $\text{proj}_{\mathcal{K}}(x_{t})=\text{proj}_{\mathcal{K}}(x_{0}+\sigma_{t}\epsilon)\approx x_{0}$, as also illustrated in Permenter & Yuan (2024).
@@ 3.3 Enhancing Sample Generation with Noise Level Correction
The trained noise level correction network $r_{\theta}$ can be integrated into various existing sampling algorithms to improve sample quality. For algorithms that include an initial sample estimate $x_{0|t}$, eq. 5, we can reformulate the one-step estimation as follows:
@@ 3.4 Constrained Sample Generation
The noise level correction method can also improve performance in constrained sample generation tasks, such as image restoration. Let $\mathcal{K}$ denote the data manifold, and let $\mathcal{C}$ represent specific constraints, such as masked pixel matching in inpainting tasks. Constrained sample generation aims to generate samples $x$ that satisfy both the manifold and constraint requirements, meaning $x\in\mathcal{K}\cap\mathcal{C}$. Similar to DDIM-NLC sampling approach in Algorithm 1, noise level correction can be incorporated into existing constrained sample generation methods, such as DDNM Wang et al. (2023), a DDIM-based image restoration algorithm. An example of this is shown in Algorithm 4.
To further enhance constrained sample generation, we propose a flexible iterative projection algorithm inspired by the alternating projection techniqueBauschke & Borwein (1996); Lewis & Malick (2008). This approach iteratively projects samples onto each constraint set to approximate a solution that lies in the intersection of $\mathcal{K}$ and $\mathcal{C}$. The iterative projection process can be expressed as follows:
@@ 4.1 Toy Experiments
As we discussed in Theorom 1, the distance to the manifold can be approximated by the noise level, as shown in Equation 13. We further evaluated the accuracy of the noise level correction by examining its proximity to the actual distance $\text{dist}_{\mathcal{K}}(\hat{x}_{t})$. Figure 4b displays the relative distance estimation bias during sampling, calculated as
@@ 4.2 Unconstrained Image Generation
We evaluate the effectiveness of noise level correction with EDM Karras et al. (2022), as it achieves state-of-the-art sampling quality with few sampling steps. Specifically, we assess the impact of NLC using both a first-order Euler ODE solver and a second-order Heun ODE solver within the EDM framework. As shown in Table 2, the proposed NLC also enhances the performance of EDM-based sampling methods. Notably, as a robust sampling technique, noise level correction improves the performance of the Heun sampler by 10% with just 13 sampling steps.
@@ 4.3 Image Restoration
In this section, we evaluate the effectiveness of noise level correction on five common image restoration tasks, 4×4$\times$4 × super-resolution (SR) using bicubic downsampling, deblurring with a Gaussian blur kernel, colorization using an average grayscale operator, compressed sensing (CS) with a Walsh-Hadamard sampling matrix at a 0.25 compression ratio, and inpainting with text masks. These experiments are conducted on the ImageNet Deng et al. (2009) and CelebA-HQ Karras et al. (2018) datasets. We compare our method with recent state-of-the-art diffusion-based image restoration methods, including ILVR Choi et al. (2021), RePaint Lugmayr et al. (2022), DDRM Kawar et al. (2022), and DDNM Wang et al. (2023). For a fair comparison, all diffusion-based methods utilize the same pretrained denoising networks with the same 100-step denoising process (100 number of inference steps), following the experimental setup in Wang et al. (2023). To evaluate sample quality, we use FID, PSNR (Peak Signal-to-Noise Ratio), and SSIM (Structural Similarity Index Measure). For colorization, where PSNR and SSIM are less effective metrics Wang et al. (2023), we additionally use a Consistency metric, denoted as "Cons" and calculated as $\|Ax_{0}-y\|_{1}$. As a baseline, we also include the inverse solution for each image restoration task, given by $\hat{x}=\textbf{A}^{\dagger}y$, which achieves zero constraint violation but lacks the data manifold information.

### v2

@@ 3.2 Noise Level Correction
For $\text{dist}_{\mathcal{K}}(x_{t})$, calculating the ground-truth distance to the manifold is generally not feasible. Instead, we approximate it using the distance between the noisy sample and its clean counterpart in the forward diffusion process. Specifically, given the noisy samples $x_{t}$ generated by the diffusion process from $x_{0}\in\mathcal{K}$ in eq. 2, we estimate the distance as $\text{dist}_{\mathcal{K}}(x_{t})\approx|x_{t}-x_{0}|$. This approximation is reasonable when the projection of $x_{t}$ onto $\mathcal{K}$ satisfies $\text{proj}_{\mathcal{K}}(x_{t})=\text{proj}_{\mathcal{K}}(x_{0}+\sigma_{t}\epsilon)\approx x_{0}$. As illustrated in Permenter & Yuan (2024), this occurs when $\sigma_{t}$ is small (due to the manifold hypothesis, the random noise $\epsilon$ is orthogonal to the manifold $\mathcal{K}$), and when $\sigma_{t}$ is much larger than the diameter of $\mathcal{K}$.
@@ 3.3 Enhancing Sample Generation with Noise Level Correction
The trained noise level correction network $r_{\theta}$ can be integrated into various existing sampling algorithms to improve sample quality. For algorithms that include an initial sample estimate $x_{0|t}$, eq. 5, we reformulate the one-step estimation as follows:
@@ 3.4 Constrained Sample Generation
The noise level correction method can also improve performance in constrained sample generation tasks, such as image restoration. Let $\mathcal{K}$ denote the data manifold, and let $\mathcal{C}$ represent specific constraints, such as masked pixel matching in inpainting tasks. Constrained sample generation aims to generate samples $x$ that satisfy both the manifold and constraint requirements, meaning $x\in\mathcal{K}\cap\mathcal{C}$. Similar to DDIM-NLC sampling approach in Algorithm 1, noise level correction can be incorporated into existing constrained sample generation methods, such as DDNM (Wang et al., 2023), a DDIM-based image restoration algorithm. An example of this is shown in Algorithm 4.
To further enhance constrained sample generation, we propose a flexible iterative projection algorithm inspired by the alternating projection technique (Bauschke & Borwein, 1996; Lewis & Malick, 2008). This approach iteratively projects samples onto each constraint set to approximate a solution that lies in the intersection of $\mathcal{K}$ and $\mathcal{C}$. The iterative projection process can be expressed as follows:
@@ 4.1 Toy Experiments
As we discussed in Theorom 1, the distance to the manifold can be approximated by the noise level, as shown in LABEL:eq:theorom_dist. We further evaluated the accuracy of the noise level correction by examining its proximity to the actual distance $\text{dist}_{\mathcal{K}}(\hat{x}_{t})$. Figure 4b displays the relative distance estimation bias during sampling, calculated as
@@ 4.2 Unconstrained Image Generation
We evaluate the effectiveness of noise level correction with EDM (Karras et al., 2022), as it achieves state-of-the-art sampling quality with few sampling steps. Specifically, we assess the impact of NLC using both a first-order Euler ODE solver and a second-order Heun ODE solver within the EDM framework. As shown in Table 2, the proposed NLC also enhances the performance of EDM-based sampling methods. Notably, as a robust sampling technique, noise level correction improves the performance of the Heun sampler by 10% with just 13 sampling steps.
@@ 4.3 Image Restoration
In this section, we evaluate the effectiveness of noise level correction on five common image restoration tasks, 4×4$\times$4 × super-resolution (SR) using bicubic downsampling, deblurring with a Gaussian blur kernel, colorization using an average grayscale operator, compressed sensing (CS) with a Walsh-Hadamard sampling matrix at a 0.25 compression ratio, and inpainting with text masks. These experiments are conducted on the ImageNet (Deng et al., 2009) and CelebA-HQ (Karras et al., 2018) datasets. We compare our method with recent state-of-the-art diffusion-based image restoration methods, including ILVR (Choi et al., 2021), RePaint plugmayr2022repaint, DDRM (Kawar et al., 2022), and DDNM (Wang et al., 2023). For a fair comparison, all diffusion-based methods utilize the same pretrained denoising networks with the same 100-step denoising process (100 number of inference steps), following the experimental setup in Wang et al. (2023). To evaluate sample quality, we use FID, PSNR (Peak Signal-to-Noise Ratio), and SSIM (Structural Similarity Index Measure). For colorization, where PSNR and SSIM are less effective metrics (Wang et al., 2023), we additionally use a Consistency metric, denoted as "Cons" and calculated as $\|Ax_{0}-y\|_{1}$. As a baseline, we also include the inverse solution for each image restoration task, given by $\hat{x}=\textbf{A}^{\dagger}y$, which achieves zero constraint violation but lacks the data manifold information.